the total cost of two calculators and four math sets is R440.00. If a calculator costs R40 more than a math set, what is the cos of each?
1
Expert's answer
2018-04-08T04:20:11-0400
Let suppose that cost of calculator and math set are c1 and c2 respectively. Then we have two equations:
2*c1 + 4*c2 = 440
c1 = c2 + 40
So
2*(c2 + 40) + 4*c2 = 440 => c2 = 360/6 = R60
and
c1 = 60 + 40 = R100